prettygirlzariah:

Tina wants to open a savings account. Bank A requires a $100 deposit to open the account and requires a minimum deposit of $50 a month. Bank B requires a $250 deposit to open the account and requires a minimum deposit of $20 a month. If Tina were to make the minimum monthly deposits in each account, enter the number of months it would take for both accounts to have the same amount of money.

BeautyQueen:

So how you solve it is 100 + 50m = 250 + 20m then you will Subtract 100 from both sides to eliminate it from the equation 50m=150 + 20m After that you Subtract 20m from both sides to eliminate it from the equation, as well 30m=150 Divide both sides by 30 to eliminate it m=5 so the answer is 5

pup:

umm ima say 5

BeautyQueen:

@pup wrote:
umm ima say 5
Ya it is the answer but we aren't allowed to just say it no more admin said we gotta explain why its the answer

pup:

cuz i say so

BeautyQueen:

@pup wrote:
cuz i say so
Chil im not finna argue with u if u give a answer show u work or they might think u wrong and when u get banned for givin answers don't go cryin have a good day ☺

BeautyQueen:

Blocked

Phantomdex:

Alright so first for Bank A, the amount after m months would be 100 + 50m. Next, we move on to Bank B, the amount after m months would be 250 + 20m. Then after that, we set the two amounts equal to each other: 100 + 50m = 250 + 20m. Lastly, we'll solve for m: 50m - 20m = 250 - 100, 30m = 150, m = 5. Which should be your answer.